Question 3 20 pts Stock XYZ has a current dividend of $3.00 The dividend is expected to grow forever at a rate of 3.00% Basedplease answer
Answer #1
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
Price = 3 * (1+0.03) / (0.04 - 0.03)
Price = 309

Dividend yield = dividend*(1+growth rate)/price = 3*(1+0.03)/309 = 1%